OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. Medals will be given, and a friend will be gained

OpenStudy (anonymous):

y = 4x - 11 y = 4x + 11 y = 4x + 7 y = 4x - 7 These are the answer choices.

Nnesha (nnesha):

slope intercept form equation \[\huge\rm y =mx + b\] m is slope b is y-intercept m=4 replace m by 4 and replace x and y by (-3,-1)

OpenStudy (anonymous):

Okay.. give me a sec to write that.

Nnesha (nnesha):

in order pair first number is x and 2nd one is y \[\huge\rm (x ,y)\]

Nnesha (nnesha):

sure :)

OpenStudy (anonymous):

Okay, (x,y) whats the next step?

Nnesha (nnesha):

replace x and y by their values and tthen solve for b

OpenStudy (anonymous):

Okay so... just to double check: y = mx + b -3 = 4 -1 + b Is this right?

OpenStudy (ribhu):

@rvc can help

Nnesha (nnesha):

x = ?? y = ??

Nnesha (nnesha):

like i said first number in order pair is x and 2nd one is y like alphabetical order x ,y ,

OpenStudy (anonymous):

x = -3 and y = -1

rvc (rvc):

@Nnesha is perfect

Nnesha (nnesha):

yes right so replace x by -3 and y by -1

Nnesha (nnesha):

thanks rvc :)

Nnesha (nnesha):

@ribhu this area is for asking question don't post comments which is not related to this post thanks use chat bars

OpenStudy (anonymous):

Okay... -1y = 4 -3x + b I wasn't sure if the x and y needed to be in there, I just put them in there for you to see.

Nnesha (nnesha):

no u have to substitute so \[\huge\rm -1 = 4(-3) +b\] and u have to write -3 in parentheses

Nnesha (nnesha):

now multiply 4 times -3

OpenStudy (anonymous):

Oh okay my bad. -12

Nnesha (nnesha):

\[\huge\rm -1 = -12 + b\] solve for b by adding 12 both side

OpenStudy (anonymous):

-1 = -12 + b +12 +12 Correct?

Nnesha (nnesha):

nope both side mean left and right side of equal sign

OpenStudy (anonymous):

-12 + -1 = -13 and than its... -12 + b, or -12 + -12?

OpenStudy (anonymous):

I have not been good at this stuff since I started. Sorry.

Nnesha (nnesha):

u have to solve for b and at right side there is already NEGATIVE 12 so add 12 both side